solve the problem
two cars leave an intersection. one car travels north; the other east. When the car traveling north has gone 9 miles, the distance between the cars was 3 miles more than the distance traveled by the car heading east. how far had the eastbound car traveled?

Answer by checkley75(3666)

HERE WE HAVE A RIGHT TRIANGLE.
9^2+X^2=12^2
81+X^2=144
X^2=144-81
X^2=63
X=SQRT63
X=7.94 ANSWER FOR THE DISTANCE THE CAR TRAVELED.